OpenStudy (anonymous):

how many solutions does the absolute value of [x-5]=4?

OpenStudy (anonymous):

It will have two. Since you need x-5=4 and x-5=-4.

OpenStudy (anonymous):

How can it be x - 5 = -4?

OpenStudy (anonymous):

9-5=4 1-5=-4(absolute value = 4) So two solutions

OpenStudy (anonymous):

oh ok.....